find the corresponding y values for x = -4,-3,-2,-1,0,1,2 if y=x2+2r-3.
answer fill in y column
x y
-4
-3
-2
-1
0
1
2
show your work here type x-squared as x ^2 unless using a superscript feature

Answer by colliefan(242)

find the corresponding y values for x = -4,-3,-2,-1,0,1,2 if y=x2+2r-3.


The equation you give must contain an error. I am guessing that it should be:
y=x^2+2x-3

In any case, you need to solve for y by substituting the different values of x into the equation and solving for y.
For the first one, if x = -4:
y=(-4)^2+2(-4)-3
y=16-8-3
y=5
and then put 5 across from -4 in the column.